I have been working on this math problem and can't figure it out and I was wondering if someone could help me? Please and Thank you! I would deeply appreciate it!
Points B and C lie on AD. Find AC if AB/BD=3/4 and AC/CD=5/6 and BD=66.

Answer by suhasbhat(3)

I assume line and points as shown A-------B-----C--------D
AB/BD = 3/4.
BD = (3/4)(AB)
But BD = 66. Therefore AB = (3/4)(BD) => AB = (3/4)*66 => AB = 99/2
Hence total length AD = AB+BD = (99/2)+ 66. => AD = 231/2.
But AD = AC+CD. => AC+CD = AD => AC + CD = 231/2.
But AC/CD = 5/6. => AC = (5/6)(CD). Now we substtute value of AC.
(5/6)CD + CD = 231/2.
Therefore CD = 693/11.
And AC = (5/6)CD.
AC = (5/6)* (693/11).
I hope you get the idea. Check the calculations.